# Restoring the Forecasting Power of Google Trends with Statistical Preprocessing

### v1

@@ 5.1 ARIMAX as a Predictive Model
The pool of keywords is obtained through Google Trends’ “related queries”, which may include terms lacking semantic relevance that cannot be manually excluded. The raw non-preprocessed dataset of individual queries is very large, while the dataset of preprocessed data is much smaller as queries are combined. To identify meaningful predictors, we calculate their correlations with the target variable and select those within the top 25th percentile for preprocessed data and 98th percentile for raw data for each state, with a maximum of 35 predictors per state. This state-specific threshold ensures that states with weaker correlations retain at least a few predictors, while states with stronger correlations include only the most relevant keywords. To prevent multicollinearity, we remove one of any pair of variables whose correlation exceeds 0.90. Correlations are computed exclusively on the training set to prevent look-ahead bias and preserve the integrity of the real-time forecasting framework.
5.2 Results
We retrospectively evaluate our approach using flu hospitalization data at both national and state (including DC) levels in the United States over the past two seasons, from October 17, 2022, to April 27, 2024. The training set extends up to October 2022. This serves as an end-to-end validation of our preprocessing pipeline, demonstrating its practical utility in real-world forecasting scenarios.
Method National State
Clustering + Denoising + Detrending (Preprocessed) 0.42 0.76
Clustering + Denoising 0.45 0.79
Clustering 0.51 0.90
Non-Preprocessed 1.54 1.22
Table 1: Relative Efficiency and Ablation Study at the US national and state levels
Table 1 displays the relative efficiencies of ARIMAX models at the US national and state levels, comparing different levels of preprocessing of exogenous variables. Since a lower MSE indicates better forecasting performance, a relative efficiency below 1 signifies an improvement over the benchmark ARIMA(1,1,1) model. Our results show that preprocessing Google Trends data enhances prediction accuracy, reducing error by 58% at the national level (RE = 0.42) and by an average of 24% at the state level (RE = 0.76). In contrast, using raw, non-preprocessed data deteriorates performance, increasing prediction error by 54% at the national level and 22% at the state level. These findings highlight the important role of preprocessing in improving predictive accuracy.
Figure 11: Summary of ratio of MSEs across all states (large outliers hidden). Lower is better.
Figure 11 extends the analysis presented in Table 1 by exploring the RE at the state level. For 88% of the states, incorporating preprocessed Google Trends data enhances model accuracy, as evidenced by the majority of the boxplot falling below 1. In contrast, the first boxplot reveals that using raw explanatory variables generally worsens performance, with only four locations showing any benefiting from unprocessed data.
We compare the forecast errors of the benchmark ARIMA model and ARIMAX with fully preprocessed variables at both national and state levels using statistical significance tests. These ensure that the observed improvements are statistically meaningful and unlikely to be due to random chance. First, we apply Student’s one-sided paired t-test (Student 1908) to test the equality of mean errors between ARIMA and ARIMAX, against the alternative hypothesis that ARIMA produces smaller errors than ARIMAX. This test is conducted on the combined data from all states, with 80 observations per location across 51 locations, resulting in a total of 4,080 observations. The overall p-value for the state-level data is $3.17\times 10^{-6}$ with a t-value of $-4.52$, well below the 0.05 significance threshold, strongly supporting rejection of the null hypothesis. This indicates that ARIMAX significantly reduces forecast error compared to ARIMA, confirming the benefit of incorporating preprocessed exogenous variables. At the national level, however, the p-value is 0.37 with a t-value of $-0.37$, showing no significant difference in the mean errors between ARIMA and ARIMAX.
To further assess forecast accuracy at the national level, we perform the Diebold-Mariano (DM) test (Diebold & Mariano 1995), a widely used method to evaluate forecast accuracy (D’Amuri & Marcucci 2017, Bantis et al. 2023). We incorporate modifications from Harvey et al. (1997), which enhance robustness in small samples, as our analysis includes only 80 observed forecast errors per location. The null hypothesis of the DM test states that two competing models have the same predictive accuracy over time. The test yields a p-value of 0.044 and a t-value of 1.73, suggesting that ARIMAX provides more stable predictive performance over time compared to ARIMA, even though it does not exhibit significantly lower average errors. The results from both tests demonstrate that the proposed end-to-end preprocessing methodology significantly improves forecasting accuracy.
@@ 6 Discussion and Conclusion
In this paper, we investigate the data generation mechanism of Google Trends and propose a comprehensive statistical preprocessing methodology to maximize its utility for forecasting. Our approach integrates hierarchical clustering to address missing values and sampling variability, dynamic smoothing splines to reduce noise, and detrending to separate short-term signals from long-term trends. Each preprocessing step is independently evaluated to validate its effectiveness in improving search data quality. We further demonstrate the practical benefits of our approach through an application to flu hospitalization prediction, showing improved accuracy when incorporating preprocessed search data into a statistical predictive model.
@@
While our method has proven effective, it is not without limitations. First, keyword selection relied on Google Trends’ “related queries” feature, which occasionally included terms unrelated to flu. Rather than manually excluding these terms, we mitigated their impact through clustering, which reduces the influence of a single irrelevant term on the aggregated search volumes, and by applying a correlation filter to exclude terms with low relevance to the target variable. Second, the clustering process occasionally grouped semantically unrelated terms, but these clusters remained meaningful in terms of their time series patterns, ensuring their utility in prediction. Third, the quality of Google Trends data varies significantly across locations, with lower internet penetration resulting in inherently poor data quality that preprocessing techniques cannot fully address. Finally, we intentionally chose a simple univariate ARIMAX model to evaluate the effectiveness of our preprocessing techniques. While more complex statistical and machine learning models could produce higher forecasting accuracy, they might obscure the direct contributions of Google Trends data. Using an interpretable model aligns with the primary focus of this paper, which is to evaluate and substantiate the effectiveness of our preprocessing methodology, rather than to develop a flu hospitalization prediction model.
@@ Appendix A Google Trends
Appendix B General ARIMA$(p,d,q)$ Model
An AutoRegressive Integrated Moving Average $\text{ARIMA}(p,d,q)$ model is used to forecast time series data. It is defined by three components, autoregression (AR), differencing (I), and moving average (MA), specified by parameters $p$, $d$ and $q$. The AR part indicates that a current value of the time series depends on $p$ of its past (or lagged) observations, while the MA part indicates it is a linear combination of $q$ of its lagged forecast errors. The I component specifies how many times ($d$) the time series is differenced to make it stationary, to stabilize the mean and variance.
The general ARIMA($p,d,q$) model can be written as [46]:
$\Delta^{d}y_{t}=\mu+\sum_{i=1}^{p}\phi_{i}\Delta^{d}y_{t-i}+\sum_{i=1}^{q}\theta_{i}\epsilon_{t-i}+\epsilon_{t}.$
The first operation $\Delta^{d}y_{t}=y_{t}-\sum_{i=1}^{d}d_{i}y_{t-i}$ represents the $d$-order differencing of the target series $y_{t}$ to achieve stationarity, with $d_{i}$ as the binomial coefficients corresponding to various orders of differencing. $\mu$ is the mean of the differenced series, $\phi_{i}$ are the coefficients for the $p$-order AR terms, which depend on past values of the differenced series $\Delta^{d}y_{t-i}$. $\theta_{i}$ are the coefficients for the $q$-order MA terms, involving past error terms $\epsilon_{t-i}$. $\epsilon_{t}$ is the current error term, assumed to follow a white noise process.
The ARIMA model is widely used for time series analysis due to its ability to model stationary and non-stationary data effectively. By integrating exogenous variables, the ARIMAX extension allows for incorporating external predictors, such as Google Trends data, to improve forecasting performance.

### v2

@@ 6.3 Modeling
The pool of keywords is obtained through Google Trends’ “related queries”, which may include terms lacking semantic relevance that cannot be manually excluded. The raw non-preprocessed dataset of individual queries is very large, therefore, we reduce it by either using topics only or by setting a zero-filter to remove keywords with too many zeros. The preprocessed dataset is much smaller, as queries are combined. To identify meaningful predictors for both raw and preprocessed data, we apply a correlation-based filter, selecting variables most correlated with the target for each location while limiting the total number of predictors to avoid overfitting. To prevent multicollinearity, we drop one variable from any pair with high correlation. Correlations are computed exclusively on the training set to prevent look-ahead bias and preserve the integrity of the real-time forecasting framework.
6.4 Forecasting Results and Statistical Significance
Forecasting performance is assessed using the Mean Squared Error (MSE) between the ground truth and predicted values, $\text{MSE}=\frac{1}{T}\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}$, over the test period t=1,,Tt=1,...,T. We then compute the Relative Efficiency (RE) as the ratio of the MSE for each model with exogenous variables to the MSE of the corresponding model without exogenous variables, which serves as the benchmark. An RE below 1 indicates improved performance over the baseline.
To evaluate statistical significance, we apply the Wilcoxon Signed-Rank Test (Wilcoxon, 1945) to the out-performance of each model. Specifically, we conduct a one-sided paired test with the null hypothesis that the median of paired differences in MSE is equal to 0 against the alternative that the median is less than 0. This nonparametric test is robust to outliers and well-suited for our 51 paired MSEs with heteroskedasticity (one per state) for each model comparison. Table 1 reports the median RE, interquartile range, and p-values aggregated across all locations, horizons, models, and sets of exogenous variables. P-values below 0.05 are indicated as * for comparisons against the baseline without exogenous information, $\dagger$ for preprocessed (clustering, denoising, and detrending) versus non-preprocessed data, and $\ddagger$ for preprocessed versus topic-only data.
Horizon Method ARIMAX SARIMAX ARGO LightGBM AdaBoost
Median (Q1, Q3) Median (Q1, Q3) Median (Q1, Q3) Median (Q1, Q3) Median (Q1, Q3)
0 Detrending 0.83∗†‡ (0.75, 0.94) 1.01†‡ (0.85, 1.17) 0.83∗† (0.69, 0.94) 0.85∗†‡ (0.76, 0.95) 0.81 (0.65, 0.98)
Denoising 0.85∗†‡ (0.75, 0.91) 1.04†‡ (0.84, 1.18) 0.80∗† (0.67, 0.92) 0.83∗†‡ (0.76, 0.90) 0.80∗† (0.66, 0.95)
Clustering 0.92∗†‡ (0.80, 0.97) 1.01†‡ (0.84, 1.23) 0.76∗† (0.55, 0.88) 0.88∗‡ (0.79, 0.97) 0.83∗† (0.68, 0.94)
Topics 1.32 (1.20, 1.63) 1.21 (1.03, 1.57) 0.78 (0.60, 0.96) 0.93 (0.84, 1.00) 0.82 (0.71, 0.97)
Non-preprocessed 1.65 (1.50, 1.87) 1.93 (1.37, 2.81) 0.94 (0.82, 1.13) 0.89 (0.81, 0.98) 0.83 (0.73, 1.03)
1 Detrending 0.80∗†‡ (0.69, 0.86) 0.81∗†‡ (0.68, 0.96) 0.72∗†‡ (0.59, 0.86) 0.71∗†‡ (0.64, 0.86) 0.62∗† (0.36, 0.74)
Denoising 0.82∗†‡ (0.70, 0.87) 0.82∗†‡ (0.70, 0.93) 0.71∗†‡ (0.59, 0.86) 0.75∗†‡ (0.60, 0.83) 0.58∗† (0.34, 0.76)
Clustering 0.82∗†‡ (0.71, 0.92) 0.80∗†‡ (0.67, 0.98) 0.66∗†‡ (0.50, 0.81) 0.80 (0.70, 0.94) 0.71 (0.49, 0.83)
Topics 1.17 (1.06, 1.31) 0.95 (0.76, 1.12) 0.92 (0.75, 1.06) 0.80 (0.73, 1.02) 0.59 (0.43, 0.86)
Non-preprocessed 1.27 (1.16, 1.43) 1.05 (0.82, 1.44) 1.05 (0.89, 1.27) 0.85 (0.71, 0.97) 0.75 (0.55, 0.91)
2 Detrending 0.83∗†‡ (0.71, 0.88) 0.72∗†‡ (0.61, 0.90) 0.73∗†‡ (0.56, 0.84) 0.75∗† (0.65, 0.88) 0.61∗† (0.44, 0.82)
Denoising 0.84∗†‡ (0.73, 0.88) 0.75∗†‡ (0.60, 0.88) 0.71∗†‡ (0.55, 0.86) 0.76∗†‡ (0.67, 0.83) 0.64∗† (0.49, 0.83)
Clustering 0.84∗†‡ (0.72, 0.92) 0.76∗†‡ (0.59, 0.85) 0.63∗†‡ (0.48, 0.82) 0.85 (0.73, 0.97) 0.75 (0.58, 0.93)
Topics 1.11 (1.01, 1.23) 0.90 (0.69, 1.12) 1.12 (1.02, 1.25) 0.83 (0.75, 0.98) 0.68 (0.45, 0.86)
Non-preprocessed 1.18 (1.04, 1.37) 0.93 (0.68, 1.20) 1.12 (1.06, 1.41) 0.86 (0.78, 1.03) 0.77 (0.60, 0.97)
3 Detrending 0.82∗†‡ (0.72, 0.90) 0.72∗†‡ (0.54, 0.89) 0.73∗†‡ (0.54, 0.87) 0.78∗† (0.66, 0.99) 0.64∗† (0.51, 0.79)
Denoising 0.84∗†‡ (0.73, 0.91) 0.72∗†‡ (0.59, 0.86) 0.74∗†‡ (0.53, 0.87) 0.81 (0.69, 1.04) 0.72∗† (0.55, 0.86)
Clustering 0.86∗†‡ (0.73, 0.94) 0.78∗†‡ (0.61, 0.89) 0.67∗†‡ (0.47, 0.80) 0.89 (0.74, 1.04) 0.78∗† (0.60, 0.90)
Topics 1.08 (0.99, 1.22) 0.87 (0.68, 1.13) 1.07 (1.00, 1.18) 0.88 (0.72, 1.02) 0.75 (0.55, 0.90)
Non-preprocessed 1.15 (1.02, 1.33) 0.88 (0.67, 1.11) 1.15 (0.98, 1.27) 0.89 (0.73, 1.03) 0.81 (0.67, 0.98)
Table 1: Median and interquartile range (Q1, Q3) of relative efficiency across all locations, horizons, models, and sets of exogenous variables. Values below 1 indicate better performance than the benchmark model without exogenous variables. Boldfaced medians correspond to the best performance relative to the baseline. p-values from the Wilcoxon Signed-Rank Test are reported as :p<0.05*:p<0.05 for models with exogenous data against baselines; $\dagger:p<0.05$ for preprocessed versus non-preprocessed data; and $\ddagger:p<0.05$ for preprocessed versus topic-only data.
Across all models and horizons, incorporating at least one preprocessing step consistently improves forecasts relative to omitting exogenous variables or using raw or topic-only data. The p-values for preprocessed versus baseline comparisons fall well below 0.05, strongly supporting rejection of the null hypothesis and confirming that preprocessing significantly reduces forecast errors. Similarly, comparisons against non-preprocessed and topic-only data show that at least one step of preprocessing outperforms raw data in most models and horizons. In contrast, using raw data degrades forecasting performance in statistical models relative to the baseline. Although the optimal level of preprocessing varies by model, these results highlight the overall benefit of applying our preprocessing pipeline to Google Trends data. Figure S2 provides boxplots illustrating variation across locations.
6.5 Comparison of Different Preprocessing Steps in Time-Varying Forecast Performance
We analyze forecasting performance separately during peak and off-season periods to assess how each preprocessing step contributes across different phases of influenza activity. Figure S3 illustrates that forecast differences between clustered, denoised, and detrended data are most pronounced during flu peaks (December 2022 and January 2024). Table 2 reports the average RE accross all locations of detrending and denoising relative to clustering alone for each forecast horizon and model. Overall, clustering by itself performs best during peak periods, as indicated by RE values above 1 for most models, whereas denoising and detrending improve accuracy during the off-season. These results suggest that denoising and detrending may overly smooth short-term fluctuations that are informative during peak outbreaks, but provide clear benefits when forecasting outside the peak season.
Horizon Season Method ARIMAX SARIMAX ARGO LightGBM AdaBoost
0 Peak Detrending 0.85 1.01 0.89 1.05 1.01
Denoising 0.89 0.97 0.89 1.05 1.01
Off Detrending 0.61 0.86 0.82 0.94 1.00
Denoising 0.57 0.80 0.74 0.94 1.00
1 Peak Detrending 1.11 1.10 0.89 1.00 1.05
Denoising 1.08 0.98 0.89 1.00 1.05
Off Detrending 0.78 0.81 0.95 1.03 0.93
Denoising 0.78 0.81 0.90 1.03 0.93
2 Peak Detrending 1.07 1.08 0.94 1.00 1.01
Denoising 1.06 0.98 0.95 1.01 1.01
Off Detrending 0.89 0.84 1.32 1.06 0.99
Denoising 0.92 0.81 1.16 1.05 0.99
3 Peak Detrending 1.01 0.99 0.97 0.9 1.00
Denoising 1.01 0.93 1.04 0.9 1.00
Off Detrending 1.02 0.91 1.70 1.04 0.88
Denoising 1.04 0.87 1.50 1.02 0.88
Table 2: Average relative efficiency between detrending/denoising and clustering across all locations, horizons, and models during peak (December–January) and off (February–November) periods over two influenza seasons. Values below 1 indicate better performance than clustering.
To examine how forecast performance evolves over time, we apply the Fluctuation Test of Giacomini and Rossi (2010) pairwise to each model’s errors using clustered, denoised, and detrended data. Unlike simple comparisons of average errors, this test assesses whether the relative performance of two methods remains stable over time. Applied to average errors across all locations for each model, the two-sided test does not reject the null hypothesis for any horizon or any pair of exogenous datasets at the confidence level 0.05. Therefore, we find no evidence that the relative performance of clustered, denoised, or detrended predictors varies over time when aggregated across locations. In other words, the three preprocessing steps perform equally well throughout the evaluation period. Figures S4-S7 display the average sample paths of the Fluctuation Test across locations for each pair of preprocessing methods, model, and forecast horizon.
@@ 7 Discussion and Conclusion
In this paper, we investigate the data generation mechanism of Google Trends and propose a comprehensive statistical preprocessing methodology to maximize its utility for forecasting. Our approach integrates hierarchical clustering to address missing values and sampling variability, dynamic smoothing splines to reduce noise, and detrending to separate short-term signals from long-term trends. Each preprocessing step is independently evaluated to validate its effectiveness in improving search data quality. We further demonstrate the practical benefits of our approach through an application to influenza hospitalization forecasting, from nowcasts up to three-week-ahead predictions, showing improved accuracy when incorporating preprocessed search data into a statistical or machine learning model.
@@
While our method has proven effective, it is not without limitations. First, keyword selection relied on Google Trends’ “related queries” feature, which occasionally included terms unrelated to flu. Rather than manually excluding these terms, we mitigated their impact through clustering, which reduces the influence of a single irrelevant term on the aggregated search volumes, and by applying a correlation filter to exclude terms with low relevance to the target variable. Second, the clustering process occasionally grouped semantically unrelated terms, but these clusters remained meaningful in terms of their time series patterns, ensuring their utility in prediction. Third, the quality of Google Trends data varies significantly across locations, with regions of lower internet penetration or smaller populations producing inherently lower quality signals that preprocessing cannot fully address. For example, California benefits from a large volume of searches, resulting in higher-quality data, while Alaska’s smaller search base produces more limited signals. This heterogeneity affects the number of meaningful clusters that can be identified in each location. In our study, we extracted a total of 56 individual terms, topics, and clusters for California, compared with only 5 for Alaska, illustrating how data sparsity potentially limits predictive ability in less populous regions.
We evaluate the effectiveness of our preprocessing techniques using a range of models, from linear statistical models, such as ARIMAX, SARIMAX, and ARGO, to nonlinear machine learning models, including LightGBM and AdaBoost. We choose interpretable models to highlight the direct contributions of Google Trends data. As each model has different specifications and levels of complexity, the appropriate level of preprocessing varies. ARIMAX and SARIMAX, which use only a few past target values, rely heavily on exogenous inputs for predictive information. For these models, denoising prevents overfitting to noise, and detrending aligns exogenous variables with the differenced target, both of which are crucial for accurate forecasts. In comparison, ARGO incorporates up to 52 weeks of past target values, which capture long-term trends, seasonal structure, and autocorrelation. In this model, exogenous variables primarily provide short-term, real-time spikes, so denoising and detrending may inadvertently remove informative fluctuations. LightGBM is a robust machine learning model that reduces the number of features through its Exclusive Feature Bundling (EFB) technique, which bundles mutually exclusive features together so that variables that contribute little to reducing prediction error are ignored (Ke et al., 2017). AdaBoost performs internal feature selection as an integral part of its procedure, since at each iteration the algorithm selects the feature that yields the best split under the current weighting (Hastie et al., 2017). As a result, these models can outperform simple baselines even with raw input data, though preprocessing further enhances performance. For instance, LightGBM, as a gradient boosting decision tree model, cannot extrapolate trends beyond the range of training values (Joseph, 2022), making detrending of exogenous predictors beneficial. AdaBoost is sensitive to noisy data, as it amplifies errors of mispredicted observations (Leistner et al., 2009), and therefore benefits from denoising. Overall, clustering is the most fundamental step of our methodology, as it overcomes privacy thresholds and reduces data sparsity by generating aggregated Google Trends search volumes. Denoising and detrending may be beneficial depending on the model. Across both statistical and machine learning approaches, at least one level of preprocessing appropriately enhances the predictive value of Google Trends data.
@@
S2 Forecasting Results
Figure S2: Ratios MSEs across all locations and horizons for different datasets, relative to the model without exogenous variables (baseline). Values below 1 indicate better performance than the baseline.
S3 Comparison of Different Preprocessing Steps in Time-Varying Forecast Performance
(a) ARIMAX
(b) SARIMAX
(c) ARGO
(d) LightGBM
(e) AdaBoost
Figure S3: Average weekly MSEs for each model across locations from October 2022 to May 2024 for clustered, denoised, and detrended data. Lower values indicate better performance. Differences are mostly observable during peaks.
Figure S4: Horizon 0 Fluctuation Test applied to average errors across all locations for each model, using a rolling window of 24 weeks. The blue line shows the path of the test statistic, with critical values indicated by thick black lines. If the path stays within the critical boundaries, we fail to reject the null hypothesis that pairs of preprocessing steps perform equally well at all time points. If the path crosses the critical values, one model outperformed the other at some point in time.
Figure S5: Horizon 1 Fluctuation Test applied to average errors across all locations for each model, using a rolling window of 24 weeks. The blue line shows the path of the test statistic, with critical values indicated by thick black lines. If the path stays within the critical boundaries, we fail to reject the null hypothesis that pairs of preprocessing steps perform equally well at all time points. If the path crosses the critical values, one model outperformed the other at some point in time.
Figure S6: Horizon 2 Fluctuation Test applied to average errors across all locations for each model, using a rolling window of 24 weeks. The blue line shows the path of the test statistic, with critical values indicated by thick black lines. If the path stays within the critical boundaries, we fail to reject the null hypothesis that pairs of preprocessing steps perform equally well at all time points. If the path crosses the critical values, one model outperformed the other at some point in time.
Figure S7: Horizon 3 Fluctuation Test applied to average errors across all locations for each model, using a rolling window of 24 weeks. The blue line shows the path of the test statistic, with critical values indicated by thick black lines. If the path stays within the critical boundaries, we fail to reject the null hypothesis that pairs of preprocessing steps perform equally well at all time points. If the path crosses the critical values, one model outperformed the other at some point in time.
S4 Model Specifications
S4.1 Denoising Models
For the SSA filter, we specify the number of lagged vectors $L$ to construct from the input time series and the number of leading components $k$ to retain. Hassani (2007) specify that $2\leq L\leq T/2$, where $T$ is the training window size of our iterative denoising algorithm, and that $L$ should be proportional to the periodicity of the time series. Since flu-related time series typically exhibit annual seasonality, we set $T=104$ weeks and $L=52$ weeks. SSA decomposes the series into a set of eigentriples representing trend, oscillatory patterns, and noise. We choose $k=3$ leading components, as the first component captures the trend and the next two form a pair that represents the annual cycle.
For the WT filter, we follow the specifications from Fenga (2020): using hard thresholding, symmetric Daubechies wavelet with 8 vanishing moments (SymmletS8), “chosen for its near symmetry and compact support”, periodic boundaries, 4-level wavelet decomposition. We optimize for the threshold value between 0.1 and 2 based on the paper’s reported optimal values.
S4.2 Forecasting Models
Traditional time series models leverage past values and external predictors and are implemented as follows. AutoRegressive Integrated Moving Average with eXogenous variables (ARIMAX) is an extension of the standard ARIMA model that incorporates external predictors. We choose a simple ARIMAX(1,1,1) model with first-order differencing, and a seasonal ARIMAX(1,1,1) model with yearly seasonal differencing. ARGO is implemented with 52 lags of the target to capture annual seasonality. These statistical models are trained on a rolling window of 104 weeks.
Tree-based machine learning approaches flexibly capture nonlinear relationships and interactions and use the following specifications. LightGBM and AdaBoost are trained on an expanding window with a minimum of four years of past data to mimic real-time forecasting and ensure sufficient training data, using all observations prior to each prediction date. For LightGBM, we include up to 52 weeks of target lags to capture seasonal dependencies and up to 4 weeks of exogenous variable lags to model short-term effects. We configure the model with the gradient-based one-side sampling (GOSS) boosting strategy, a maximum tree depth of 5, 300 boosting iterations, and a learning rate of 0.5. For AdaBoost, we restrict the input to 1–4 weeks of target and exogenous lags to emphasize recent dynamics, and use decision trees of maximum depth 3 as base learners with 500 boosting iterations and a learning rate of 0.5. These configurations are designed to balance predictive flexibility with efficiency, limiting overfitting while maintaining computational feasibility in the rolling forecasting setup that requires frequent retraining.